OpenStudy (anonymous):

How do you express sin 4theta in terms of theta

OpenStudy (anonymous):

You need two identities:\[\sin2x=2(\sin x)( \cos x)~~~and~~~\cos2x=\cos^2x-\sin^2x\]Once we have these, it's pretty direct.\[\sin 4x=2(\sin 2x)(\cos 2x)=4(sinx)(cosx)(\cos^2x-\sin^2x)\]

OpenStudy (anonymous):

The identities come from formulas for sin of a sum and cos of a sum.

OpenStudy (anonymous):

Do math every day.

OpenStudy (anonymous):

Thanks, expressing in terms of theta confused me